Problem 1: You are considering investing in a 10-year bond issued by NewEnergy Inc. This bond has $1000 face value, 4% coupon rate.
The bond pays coupons semi-annually and is currently selling at $920. The bond can be called at a $1,040 in 3 years.
1.a. (10 points): If your required rate of return if 6% for bonds in this risk class, what is the maximum price you should pay for this bond? (Use PV function)
Coupon rate=
Required return= N=
Years to maturity= I/Y=
Payment frequency= FV=
Face Value= PMT=
Call price= PMT Type=
Current market price= PV=
the maximum price you should pay for this bond:
1.b. (10 points) What is the current yield of this bond?
Current yield =
Answer #1

1]

1.a]

The bond is trading at a discount, which means that the interest rates have increased since the bond was issued. Hence, the bond is not likely to be called. The maximum price of bond is calculated assuming that the bond will not be called, and will be held until maturity.

I/Y = 3% (Semiannual required return = annual required return / 2 = 6% / 2 = 3%)

N = 20 (10 years remaining until maturity with 2 semiannual coupon payments each year)

PMT = 20 (semiannual coupon payment = face value * coupon rate / 2 = $1000 * 4% / 2 = $20)

FV = 1000 (face value of bond receivable at maturity)

PV is calculated to be 851.23.

The maximum price of bond is $851.23.

1.b]

Current yield = annual coupon payment / bond price.

annual coupon payment = face value * coupon rate = $1000 * 4% = $40.

Current yield = $40 / $920 = 4.35%